# Cosmology with Binary Neutron Stars: Does the Redshift Evolution of the Mass Function Matter?

### v1

@@ Abstract
Next-generation gravitational wave detectors are expected to detect millions of compact binary mergers across cosmological distances. The features of the mass distribution of these mergers, combined with gravitational wave distance measurements, will enable precise cosmological inferences, even without the need for electromagnetic counterparts. However, achieving accurate results requires modeling the mass spectrum, particularly considering possible redshift evolution. Binary neutron star (BNS) mergers are thought to be less influenced by changes in metallicity compared to binary black holes (BBH) or neutron star-black hole (NSBH) mergers. This stability in their mass spectrum over cosmic time reduces the chances of introducing biases in cosmological parameters caused by redshift evolution. In this study, we use the population synthesis code COMPAS to generate astrophysically motivated catalogs of BNS mergers and explore whether assuming a non-evolving BNS mass distribution with redshift could introduce biases in cosmological parameter inference. Our findings demonstrate that, despite large variations in the BNS mass distribution across binary physics assumptions and initial conditions in COMPAS, the mass function remains redshift-independent, allowing a 2% unbiased constraint on the Hubble constant — sufficient to address the Hubble tension. Additionally, we show that in the fiducial COMPAS setup, the bias from a non-evolving BNS mass model is less than $0.5\%$ for the Hubble parameter measured at redshift $0.4$. These results establish BNS mergers as strong candidates for spectral siren cosmology in the era of next-generation gravitational wave detectors.
@@ 1 Introduction
Estimating redshift only from GWs is not immune to systematic uncertainties. One of the most promising techniques rely on the fact that GWs redshift mass scales (for an alternative approach using neutron star equation of state to infer redshift, see Messenger & Read (2012)). Identifying features, such as peaks or dips, in the mass distribution at different luminosity distances can estimate redshifts from a catalog of CBCs (Chernoff & Finn, 1993; Taylor et al., 2012; Taylor & Gair, 2012; Farr et al., 2019), a method known as the spectral-siren method (Ezquiaga & Holz, 2022):
@@
To address bias from the redshift evolution of the mass spectrum, we can model the fully correlated population of masses and redshifts, denoted as $\pi(M_{c},q,z)$. Here, $M_{c}$ is the source-frame chirp mass, $q$ is the mass ratio, and we exclude spins since we are focusing on binary neutron star mergers, where spins are expected to be small (see section 4.3 of Chattopadhyay et al. (2020), this is also consistent with galactic observations, see Zhu et al. (2018) and Zhu & Ashton (2020)). However, there is no universally accepted astrophysics-driven model for $\pi(M_{c},q,z)$ due to significant uncertainties in the astrophysical processes and quantities that govern the population. Without a model for redshift evolution, analyses often default to a simplifying assumption, treating mass models as redshift-independent, where $\pi(M_{c},q,z)=\pi(M_{c},q)\pi(z)$, where $\pi(M_{c},q)$ and $\pi(z)$ represent the uncorrelated population distributions of $M_{c}$ and $q$, and $z$, respectively.
Several astrophysical indicators suggest that the mass distribution of BNS evolves much less with redshift compared to the binary black hole (BBH) and neutron star-black hole (NSBH) merger populations. In this paper, we aim to test whether assuming redshift independence in the BNS mass population introduces biases in the inferred cosmological parameters. It is, at present, impossible to observationally assess the redshift evolution of the BNS population because GWTC-3 contains only the two BNS mergers GW170817 and GW190425 (Abbott et al., 2023b). The limited sensitivity, and associated small search volumes, for the present generation of detectors will make measurements of redshift evolution difficult even with larger catalogs.
@@
    Is a redshift-evolving mass function for BNS mergers necessary to address the Hubble tension?
@@
    At which redshift can we best measure the Hubble parameter, and can a redshift-unevolving mass function provide an unbiased, sub-percent estimate of it there?
To explore these questions, we generated multiple astrophysically motivated BNS catalogs using the population synthesis code COMPAS (Stevenson et al., 2017; Vigna-Gómez et al., 2018; Broekgaarden et al., 2019; Riley et al., 2022), varying the physical assumptions and initial parameters. For each catalog, we simulated observations with xG observatories and estimated cosmological parameters using models of the mass function that do not evolve with redshift to assess whether the true values were accurately recovered.
We want to clarify that in this paper we do not infer population together with cosmology. Our focus is different: we test whether cosmological parameters inferred from the injected population remain consistent with the assumption of a non-evolving mass distribution. The BNS mass distribution varies significantly with binary physics (see, for example, Fig. 1 in van Son et al., in prep), and there is currently no astrophysically motivated parametric model that covers all BNS mass distributions produced by population synthesis frameworks. Additionally, mass measurements from both electromagnetic and gravitational wave observations are still too sparse to confirm which synthesized mass distribution is correct (see Appendix A). Here, we investigate whether, despite variations in the BNS mass distribution under different initial conditions and binary physics choices in COMPAS, it remains redshift-independent to address the Hubble tension.
@@ 2 Redshift Evolution of Features in Mass Distribution
Firstly, there are fewer formation channels that contribute to the population of BNS mergers. While BBH mergers are anticipated to form through a wide range of channels including dynamical interactions, homogeneous evolution, and classical isolated binary evolution (e.g., Mandel & Farmer, 2022), the formation channels for BNS mergers are much more constrained. Specifically, BNSs are expected to form primarily (if not exclusively) from isolated binary channels involving at least one common envelope event (see e.g., Fig. 1 from Wagg et al. 2022 and, Fig. 6 from Iorio et al. 2023). The reason behind this is that BNS mergers are not expected to form through the ‘stable mass transfer channel’ due to limitations imposed by the critical mass ratio required for this channel (e.g., van Son et al., 2022b; Picco et al., 2024). This implies that we primarily need to focus on the redshift dependence and delay-time distribution of the ‘common envelope channel,’ which is expected to produce relatively short delay times peaking around 1 Gyr (e.g., Fishbach & van Son, 2023), leading to a BNS merger rate that closely tracks the overall star formation rate.
Second, BNS formation (or the yield per unit of star formation) has been shown to be much less sensitive to, or even unaffected by, the formation metallicity. This is in stark contrast to BBH formation, that has repeatedly been shown to drop at high metallicities. This behavior is consistently observed across various population synthesis models with a very broad variation of physics assumptions (including BPASS, MOBSE, StarTrack, COMPAS, SEVN Eldridge & Stanway, 2016; Giacobbo & Mapelli, 2018; Giacobbo et al., 2018; Klencki et al., 2018; Neijssel et al., 2019; Santoliquido et al., 2021; Broekgaarden et al., 2022; Iorio et al., 2023; Fishbach & van Son, 2023). DCO formation through the CE channel is influenced by metallicity-dependent winds, which reduce the progenitor core masses. For BBHs, lower-mass BHs receive larger natal kicks because these kicks are assumed to scale inversely with mass. Hence this leads to a loss of potential BBHs at high metallicities. In contrast, neutron star progenitors are not assumed to experience mass-scaled natal kicks, and thus BNS formation is not sensitive to metallicity in the same way (see van Son et al., in prep for a more in depth discussion of the origin and robustness of this result). This lack of metallicity dependence suggests that the BNS mass distribution will be less impacted by the changing metallicity distribution of star formation across redshifts (Chruślińska, 2024).
@@
Figure 1: The absolute value of the Pearson correlation coefficient (linear correlation) between mass and redshift for the primary (blue) and secondary (red) neutron star masses across population synthesis variations. Simulation variant names are detailed in the text. The absolute linear correlation remains below 0.02 for all variations. Among these, the Mandel-Müller model shows the highest correlation between primary mass and redshift. Code to generate these plots is available here.
To verify this expectation, we explore variations in uncertain astrophysical parameters that govern BNS mergers, as shown in Figure 1. We focus on physics variations that are expected to have a potentially significant impact on the population of BNS mergers. The population synthesis variations considered are:
@@
    Mass transfer stability parameter for Hertzsprung gap donors, $\zeta_{\text{HG}}=5,~{}5.5,~{}6$ (see Riley et al., 2022, for more details).
@@
    Width of the Maxwellian distribution of the kick velocity for electron capture supernovae, $\sigma_{\text{ECSN}}=10,~{}200$ (km/s).
As outlined in the COMPAS methods paper (Riley et al., 2022), the fiducial model assumes CE $\alpha$=1, $\zeta_{\text{HG}}=6.5$, and $\sigma_{\text{ECSN}}=30$ (km/s). It adopts the ‘Delayed’ prescriptions from Fryer et al. (2012) for the remnant masses and natal kicks, and has an accretion efficiency $\beta$, that is limited by the thermal timescale of the donor star. To calculate the star formation rate density (SFRD) as a function of redshift and metallicity, we use the cosmic integrator from the COMPAS suite (Neijssel et al., 2019), following the parameters outlined in van Son et al. (2023).
The absolute value of the mass-redshift correlation can increase up to $10$ times compared to the fiducial population (see Figure 1), with the Mandel & Müller (2020) kick prescription. However, this correlation remains below $0.02$ for all variations explored, and including non-linear correlations does not impact the results (see full contour plots here). The low correlation observed between mass and redshift across these models supports our assumption that the mass distribution is nearly independent of redshift.
@@ 3 Spectral Siren Cosmology with an Evolving BNS Mass Spectrum
To test whether assuming a redshift-independent BNS mass population biases our cosmological inference, we conduct an injection-recovery campaign by generating mock GW data for BNS mergers. The mock GW data is generated in Cosmic Explorer sensitivity (Evans et al., 2021), with the injected population assumed to be fiducial and Mandel-Müller. We select the Mandel-Müller model because it exhibits the highest mass-redshift correlation (see Fig. 1). If this model still allows for unbiased cosmological parameter inference, then the other models should also be safe.
Figure 2: Injection-recovery results for the fiducial model (top) and the Mandel & Müller (2020) remnant-mass and natal kick prescription (bottom). Left: Kernel density estimates of the $m_{1}$ $z$ and $m_{2}$ $z$ populations from COMPAS simulations. In both models, peaks and dips remain stationary with redshift, changing only in absolute merger rate based on star formation, suggesting that the BNS mass function is roughly redshift-independent. Right: Inferred matter density and Hubble constant values for both correlated (blue) and uncorrelated (green) populations. Our framework accurately recovers cosmological parameters using both evolving and non-evolving mass models, based on mock data from $800$ binary neutron stars (about 6 days of xG observations at the median BNS merger rate (Abbott et al., 2023b)). For both the fiducial and Mandel & Müller (2020) models, we achieve $\sim 2\%$ constraints on the Hubble constant, which is sufficient to address the Hubble tension. Code to generate these plots is available here and here.
@@
Our population synthesis simulations provide discrete three-dimensional samples of masses and redshifts. We first fit the population with a three-dimensional Gaussian Mixture Model in chirp mass, mass ratio, and redshift space, and then resample from it. We apply a signal-to-noise ratio (SNR) cut of 8 and adopt the Planck 2018 cosmology (Aghanim et al., 2020) to convert this into the detected population. Additionally, we incorporate parameter estimation (PE) uncertainties from GWs to obtain samples of detector-frame masses and luminosity distances for each BNS binary. See Appendix C for details about the mock data generation.
We validate our cosmological inference framework using the injected fully correlated population of masses and redshift, represented by $\pi(M_{c},q,z)$. Subsequently, we test the assumption of an uncorrelated population, where $\pi(M_{c},q,z)\approx\pi(M_{c},q)\pi(z)$. To model the uncorrelated mass and redshift populations, $\pi(M_{c},q)$ $\pi(z)$, respectively, we fit them individually using Gaussian Mixture Models with marginalized samples of $M_{c}$, $q$ and $z$ obtained from the injected distribution. Here and throughout our models are assuming that the population distribution is perfectly known. This produces maximally precise cosmological inferences and is therefore more sensitive to biases from unmodeled mass-redshift correlation. A realistic analysis that simultaneously fits the population distribution and cosmology will be even less sensitive to these biases at the cost of additional uncertainty in the cosmological parameters. For further details regarding the inference framework, please refer to Appendix D.
Fig. 2 presents the injection-recovery results for mock data generated under fiducial assumptions and the Mandel-Müller prescription for 800 detected BNS mergers. For both the fiducial and Mandel-Müller data, the joint posterior distributions of the Hubble constant ($H_{0}$) and matter density parameter ($\Omega_{M}$), derived using the assumption of non-evolving mass spectra (uncorrelated case), align with the results obtained when considering the injected population (which may exhibit mass-redshift correlations, i.e., the correlated case). Both methods accurately recover the injected Planck 2018 values with $\sim 2\%$ precision in $H_{0}$, suggesting that a redshift-evolving BNS mass model is not required to address the Hubble tension across all astrophysical variations considered in Section 2.
In Fig. 3, we extend the injection-recovery campaign using fiducial mock data to cover up to one month of observation time with cosmic explorer. No bias is observed when using a non-evolving population model for inference, achieving a $1.14\%$ constraint on $H_{0}$. For $3,000$ BNS mergers in the fiducial universe, the Hubble parameter, $H(z)$, is best constrained at $z=0.4$, with less than $0.5\%$ bias from any redshift evolution in the mass function. This suggests that we can make sub-percent level conclusions about the cosmic history of our universe if it follows the fiducial setup of COMPAS.
@@ 4 Conclusion and Future Work
In this paper, we have evaluated BNS populations as spectral sirens, specifically examining the bias introduced by assuming a redshift-uncorrelated population, $\pi(M_{c},q)\pi(z)$, compared to the injected population, $\pi(M_{c},q,z)$. Our analysis shows that using a non-evolving mass model allows us to recover the injected Hubble constant and matter density parameter accurately (see Fig. 2), across a range of realistic population synthesis models, including a fairly extreme example of mass-redshift correlation in the Mandel-Müller kick prescription. Our results indicate that a redshift-evolving mass model is not required to achieve a $2\%$ constraint on the Hubble constant under reasonable physical assumptions about the evolution of the BNS mass function with redshift. This level of precision in $H_{0}$ is sufficient to make a significant contribution toward resolving the Hubble tension. Extending our injection-recovery campaign to one month of Cosmic Explorer data, we have found that a redshift-evolving mass model is not required to achieve sub-percent level precision in the Hubble constant provided BNS mergers in our universe are similar to those of the fiducial setup of COMPAS.
In this study, we have focused on variations of binary physics parameters that are both widely regarded as uncertain, and that we expect could significantly affect the metallicity dependence of the BNS mass distribution. However, our analysis is by no means exhaustive. The BNS mass distribution may still depend on metallicity, and therefore evolve with redshift, if parameters like the neutron star natal kick distribution or common envelope physics vary intrinsically with metallicity. In future work, we plan to explore such variations to determine whether the BNS distribution can be made redshift-dependent. Nonetheless, we anticipate that the BNS mass distribution is far less prone to redshift evolution than the BBH/NSBH mass distribution (see section 2).

### v2

# Cosmology with Binary Neutron Stars: Does Mass-Redshift Correlation Matter?

@@ Abstract
Next-generation gravitational wave detectors are expected to detect millions of compact binary mergers across cosmological distances. The features of the mass distribution of these mergers, combined with gravitational wave distance measurements, will enable precise cosmological inferences, even without the need for electromagnetic counterparts. However, achieving accurate results requires modeling the mass spectrum, particularly considering possible redshift evolution. Binary neutron star (BNS) mergers are thought to be less influenced by changes in metallicity compared to binary black holes (BBH) or neutron star-black hole (NSBH) mergers. This stability in their mass spectrum over cosmic time reduces the chances of introducing biases in cosmological parameters caused by redshift evolution. In this study, we use the population synthesis code COMPAS to generate astrophysically motivated catalogs of BNS mergers and explore whether assuming a non-evolving BNS mass distribution with redshift could introduce biases in cosmological parameter inference. . Additionally, we show that in the fiducial COMPAS setup, the bias from a non-evolving BNS mass model is less than $0.5\%$ for the Hubble parameter measured at redshift $0.4$. These results establish BNS mergers as strong candidates for spectral siren cosmology in the era of next-generation gravitational wave detectors.
@@ 1 Introduction
Estimating redshift only from GWs is one of the most promising techniques rely on the fact that GWs redshift mass scales (for an alternative approach using neutron star equation of state to infer redshift, see Messenger & Read (2012)). Identifying features, such as peaks or dips, in the mass distribution at different luminosity distances can estimate redshifts from a catalog of CBCs (Chernoff & Finn, 1993; Taylor et al., 2012; Taylor & Gair, 2012; Farr et al., 2019), a method known as the spectral-siren method (Ezquiaga & Holz, 2022):
@@
It is, at present, impossible to observationally assess the redshift evolution of the BNS population because GWTC-3 contains only the two BNS mergers GW170817 and GW190425 (Abbott et al., 2023b). The limited sensitivity, and associated small search volumes, for the present generation of detectors will make measurements of redshift evolution difficult even with larger catalogs.
@@
To explore these questions, we generated multiple astrophysically motivated BNS catalogs using the population synthesis code COMPAS (Stevenson et al., 2017; Vigna-Gómez et al., 2018; Broekgaarden et al., 2019; Riley et al., 2022), varying the physical assumptions and initial parameters. we simulate observations with xG observatories and estimate cosmological parameters using models of the mass function that with redshift to assess whether the true values are accurately recovered.
We want to clarify that in this paper we do not infer population together with cosmology. Our focus is different: we test whether cosmological parameters inferred from the injected population remain consistent with the assumption of a non-evolving mass distribution. The BNS mass distribution varies significantly with binary physics (see, for example, Fig. 1 in van Son et al. (2024)), and there is currently no astrophysically motivated parametric model that covers all BNS mass distributions produced by population synthesis frameworks. Additionally, mass measurements from both electromagnetic and gravitational wave observations are still too sparse to confirm which synthesized mass distribution is correct (see Appendix A). Here, we investigate whether, despite variations in the BNS mass distribution under different initial conditions and binary physics choices in COMPAS, it remains sufficiently redshift-independent to address the Hubble tension. Ultimately,
@@ 2 Redshift Evolution of Features in Mass Distribution
Firstly, there are fewer formation channels that contribute to the population of BNS mergers. While BBH mergers are anticipated to form through a wide range of channels including dynamical interactions, homogeneous evolution, and classical isolated binary evolution (e.g., Mandel & Farmer, 2022), BNSs are expected to form primarily (if not exclusively) from isolated binary channels involving at least one common envelope event (see e.g., Fig. 1 from Wagg et al. 2022 and, Fig. 6 from Iorio et al. 2023). The reason behind this is that BNS mergers are not expected to form through the ‘stable mass transfer channel’ due to limitations imposed by the critical mass ratio this channel (for details, see van Son et al., 2022b; Picco et al., 2024). This implies that we primarily need to focus on the redshift dependence and delay-time distribution of the ‘common envelope channel,’ which is expected to produce relatively short delay times peaking around 1 Gyr (e.g., Fishbach & van Son, 2023) leading to a BNS merger rate that closely tracks the overall star formation rate.
Second, BNS formation (or the yield) has been shown to be much less sensitive to, or even unaffected by the formation metallicity. This is in stark contrast to BBH formation, where has repeatedly been shown to drop at high metallicities. This behavior is consistently observed across various population synthesis models with a very broad variation of physics assumptions (including BPASS, MOBSE, StarTrack, COMPAS, SEVN Eldridge & Stanway, 2016; Giacobbo & Mapelli, 2018; Giacobbo et al., 2018; Klencki et al., 2018; Neijssel et al., 2019; Santoliquido et al., 2021; Broekgaarden et al., 2022; Iorio et al., 2023; Fishbach & van Son, 2023). DCO formation through the CE channel is influenced by metallicity-dependent winds, which reduce the progenitor core masses . For BBHs, lower-mass BHs receive larger natal kicks because these kicks are assumed to scale inversely with mass . Hence this leads to a loss of potential BBHs at high metallicities. In contrast, neutron star progenitors are not assumed to experience mass-scaled natal kicks, and thus BNS formation is not sensitive to metallicity in the same way (see for a more in depth discussion of the origin and robustness of this result). This lack of metallicity dependence suggests that the BNS mass distribution will be less impacted by the changing metallicity distribution of star formation across redshifts (Chruślińska, 2024).
@@
The population synthesis variations considered are:
@@
    $\zeta_{\text{HG}}=5,~{}5.5,~{}6$ (see Riley et al., 2022, for more details).
@@
    $\sigma_{\text{ECSN}}=10,~{}200$ (km/s).
To calculate the star formation rate density (SFRD) as a function of redshift and metallicity, we use the cosmic integrator from the COMPAS suite (Neijssel et al., 2019), following the parameters outlined in van Son et al. (2023).
2.1 Population Synthesis Variations Not Considered that could Potentially Show Mass-Redshift Correlation
  1. 1.
  2. 2.
@@ 3 Spectral Siren Cosmology with an Evolving BNS Mass Spectrum
To test whether assuming a redshift-independent BNS mass population biases our cosmological inference, we conduct an injection-recovery campaign by generating mock GW data for BNS mergers. The mock GW data is generated in Cosmic Explorer sensitivity (Evans et al., 2021), with the injected population assumed to be fiducial and Mandel-Müller. We select the Mandel-Müller model because it exhibits the highest mass-redshift correlation (see . If this model still allows for unbiased cosmological parameter inference, then the other models should also be safe.
Figure 1: Injection-recovery results for the fiducial model (top) and the Mandel & Müller (2020) remnant-mass and natal kick prescription (bottom). Left: Kernel density estimates of the $m_{1}$ $z$ and $m_{2}$ $z$ populations from COMPAS simulations. In both models, peaks and dips remain stationary with redshift, changing only in absolute merger rate based on star formation, suggesting that the BNS mass function is roughly redshift-independent. Right: Inferred matter density and Hubble constant values for both correlated (blue) and uncorrelated (green) populations. Our framework accurately recovers cosmological parameters using both evolving and non-evolving mass models, based on mock data from $800$ binary neutron stars (about 6 days of xG observations at the median BNS merger rate (Abbott et al., 2023b)). For both the fiducial and Mandel & Müller (2020) models, we achieve $\sim 2\%$ constraints on the Hubble constant, sufficient to address the Hubble tension. Code to generate these plots is available here and here.
@@
Our population synthesis simulations provide discrete three-dimensional samples of masses and redshifts. We first fit the population with a three-dimensional Gaussian Mixture Model in chirp mass, mass ratio, and redshift space, and then resample from it. We apply a signal-to-noise ratio (SNR) cut of 8 and adopt the Planck 2018 cosmology (Aghanim et al., 2020) to convert this into the detected population. Additionally, we incorporate parameter estimation (PE) uncertainties from GWs to obtain samples of detector-frame masses and luminosity distances for each BNS binary.
We validate our cosmological inference framework using the injected fully correlated population of masses and redshift, represented by $\pi(M_{c},q,z)$. Subsequently, we test the assumption of an uncorrelated population, where $\pi(M_{c},q,z)\approx\pi(M_{c},q)\pi(z)$. . Here and throughout our models are assuming that the population distribution is perfectly known. This produces maximally precise cosmological inferences and is therefore more sensitive to biases from unmodeled mass-redshift correlation. A realistic analysis that simultaneously fits the population distribution and cosmology will be even less sensitive to these biases at the cost of additional uncertainty in the cosmological parameters. For further details regarding the inference framework, please refer to Appendix D.
Fig. 1 presents the injection-recovery results for mock data generated under fiducial assumptions and the Mandel-Müller prescription for 800 detected BNS mergers. For both the fiducial and Mandel-Müller data, the joint posterior distributions of the Hubble constant ($H_{0}$) and matter density parameter ($\Omega_{M}$), derived using the assumption of non-evolving mass spectra (uncorrelated case), align with the results obtained when considering the injected population (which may exhibit mass-redshift correlations, i.e., the correlated case). Both methods accurately recover the injected Planck 2018 values with $\sim 2\%$ precision in $H_{0}$, suggesting that .
In Fig. 2, we extend the injection-recovery campaign using fiducial mock data to cover up to one month of observation time with cosmic explorer. No bias is observed when using a population model for inference, achieving a $1.14\%$ constraint on $H_{0}$. For $3,000$ BNS mergers in the fiducial universe, the Hubble parameter, $H(z)$, is best constrained at $z=0.4$, with less than $0.5\%$ bias from any . This suggests that we can make sub-percent level conclusions about the cosmic history of our universe if it follows the fiducial setup of COMPAS.
@@ 4 Conclusion and Future Work
In this paper, we have evaluated BNS populations as spectral sirens, specifically examining the bias introduced by assuming a redshift-uncorrelated population, $\pi(M_{c},q)\pi(z)$, compared to the injected population, $\pi(M_{c},q,z)$. Our analysis shows that using a mass model allows us to recover the injected Hubble constant and matter density parameter accurately (see Fig. 1), across a range of realistic population synthesis models, including a fairly extreme example of mass-redshift correlation in the Mandel-Müller kick prescription. Our results indicate that a redshift- mass model is not required to achieve a $2\%$ constraint on the Hubble constant under reasonable physical assumptions about the evolution of the BNS mass function with redshift. This level of precision in $H_{0}$ is sufficient to make a significant contribution toward resolving the Hubble tension. Extending our injection-recovery campaign to one month of Cosmic Explorer data, we have found that a redshift- mass model is not required to achieve sub-percent level precision in the Hubble constant provided BNS mergers in our universe are similar to those of the fiducial setup of COMPAS.
In this study, we have focused on variations of binary physics parameters that are both widely regarded as uncertain, and that we expect could significantly affect the metallicity dependence of the BNS mass distribution. However, our analysis is by no means exhaustive. The BNS mass distribution may still depend on metallicity, and therefore evolve with redshift, if parameters like the neutron star natal kick distribution or common envelope physics vary intrinsically with metallicity Nonetheless, we anticipate that the BNS mass distribution is far less prone to redshift evolution than the BBH/NSBH mass distribution (see section 2).